# Towards Better Visualizing the Decision Basis of Networks via Unfold and Conquer Attribution Guidance

### v1

@@ Abstract
Revealing the transparency of Deep Neural Networks (DNNs) has been widely studied to describe the decision mechanisms of network inner structures. In this paper, we propose a novel post-hoc framework, Unfold and Conquer Attribution Guidance (UCAG), which enhances the explainability of network decision by spatially scrutinizing the input features with respect to the model confidence. Inspired by the phenomenon that sub-regions are disregarded in the original attribution maps, UCAG sequentially disambiguates the confidence of slices within the image along the spatial axis, leading to providing a rich and clear interpretation. Therefore, it is possible to enhance the representation ability of attribution by preserving the detailed descriptions of assistant input features, which are commonly overwhelmed by the main meaningful regions. We conduct numerous evaluations to validate the performance in several metrics: i) deletion and insertion, ii) (energy–based) pointing games, and iii) positive and negative density maps. Experimental results, including qualitative comparisons, demonstrate that our method outperforms the existing methods with the nature of clear and detailed explanations and applicability.
@@ Introduction
Although Deep Neural Networks (DNNs) show remarkable performance in various computer science tasks due to their applicability, the issues of complexity and transparency derived from the non-linearity of the network hinder the clear interpretation of the decision basis. Visual explanations of network decision alleviate these problems and provide analytical guidance for model development with the intuitive explanations. Inspired by this, several approaches have been introduced, such as relevance or gradient–based generation, to describe the image regions that the network currently references to make decisions.
@@
Figure 1: The qualitative results of our approach. First and second rows represent the existing visualization methods and ours, respectively. Our approach enhances representation ability of explanation maps, applicable to any explanation techniques.
@@
The approaches behind the relevance score generate the explanation based on the decomposition methods, Layer-wise Relevance Propagation (LRP) (bach2015pixel). LRP constructs the propagation rule according to the predicted scores of the specific class. Despite the nature of employing relevance score is robust to the shattered gradient problem and success in producing the input-level explanation map, the class-discriminativeness within their representation is still limited. Recently, although Attribution Guided Factorization (AGF) enhances the class discriminability of the explanation by integrating both gradient/relevance–based methodologies, it still provides the biased attribution toward the class.
@@
GradCAM (selvaraju2017grad) computes the coefficients for each map which are averaged over the gradients of the activation maps and generates a class activation map by linearly combining the maps with the corresponding coefficients to provide a visual explanation. Although the explanation map generated from GradCAM successfully provides the class-discriminative explanation, the provided explanations are still coarse and contain numerous non-saliency regions due to the mechanism of utilizing the extraction source map from the deepest layer, which keeps the lowest resolution activation map.
The compressive nature of DNNs limits the representation ability of relevance/gradient–based approaches. As the features in deep layer of the network holds the abbreviated information, restricts the propagated regions to represent abundant saliencies. Thus, the resulting saliency maps produced from the gradient or relevance approach are represented with limited quality, with obscured class-discriminative regions.
@@
In this paper, inspired by the divide-and-conquer philosophy, we propose the Unfold-and-Conquer Attribution Guidance (UCAG) framework to represent better regions that have contributed to the decision-making process. UCAG improves the quality of explanations by allowing an attribution approach to scrutinize a set of subregions enumerated from a spatial perspective. To fulfill this, UCAG is divided into three sections: spatial unfoldment, partial saliency generation, and conquer with geometrical aggregation. First, the unfoldment procedure provides a series of patches delivered by deploying the original data from a spatial perspective. The attribute method independently computes the partial explanation map. During this procedure, UCAG accepts any specific type of methodology as an explanation module, including attribution mechanisms and model types. Finally, the conquering procedure aggregates the partial saliency maps to produce the final explanatory map. Therefore, UCAG encourages an attribution method that scrutinizes an image through the spatial axis.
The key contributions of this paper is summarized as:
@@
  • We propose a new post-hoc framework: Unfold and Conquer Attribution Guidence (UCAG) to improve the explanability of attribution methods by spatially scrutinizing the input data. Spatial scrutinization and integration provide multiple partial semantic views, leading to exploring the obscured saliencies due to the compressing nature of DNNs.
@@
  • Post-hoc manner of UCAG has the advantage of easily applicable to various relevance/gradient–based approaches and DNN models. We present the applicability of UCAG that successfully provides a function of post-hoc refinement with the intuitive and detailed descriptions of explanation.
@@
  • We validate the quality of saliency map in density, localization, and causality perspectives. We conduct the verified experiments: i) positive and negative map densities (jalwana2021cameras), ii) energy–based and original pointing game (wang2020score), and iii) insertion and deletion (petsiuk2018rise) to confirm the perspectives, respectively. The results demonstrate that our method outperforms the existing explaining methods with a sizable gap in terms of indicating the better points in the input correlated with the output decision.
@@
Starting with CAM, numerous approaches introduce strategies that provide a clear explanation based on activation maps and their coefficients. We divided the flow of research into two sections: the gradient/relevance–based approach. Gradient–based approaches (selvaraju2017grad; chattopadhay2018grad; fu2020axiom) construct the coefficients by employing the gradients corresponding to the activation map. GradCAM alleviates the limitations of CAM that require a global average pooling and re-training process. They construct an explanatory map by linearly combining the activation map and the coefficients with the coefficients computed by taking the average of the gradients corresponding to the activation map. XGradCAM (fu2020axiom) introduces the way to generate the class activation map while obeying two theoretical axioms: sensitivity and conservation. WGradCAM (zhang2021novel) enhances the representation ability of the explanation map in terms of localization.
@@
Several approaches (jiang2021layercam; jalwana2021cameras; zhang2021novel) turned the literature into amplifying the quality of activation maps to produce a more precise saliency map. LayerCAM (jiang2021layercam) aggregates the information from earlier layers and introduces the way of deciding the coefficients of each source. CAMERAS (jalwana2021cameras) fuses the multi-resolution images to generate a high-resolution explanation while obeying the model fidelity. SigCAM (zhang2021novel) combines perturbation/gradient–based approaches to produce more sparse and precise explanations using learnable external parameters.
In contrast to the representation ability within the input level being limited in CAM–based approaches, relevance–based approaches (nam2020relative; nam2021interpreting; gur2021visualization) successfully produce the granular explanation based on the LRP (bach2015pixel). LRP employs the propagation rule according to the scores predicted by the model with the input image. As the variants of the approach, Relative Attributing Propagation (RAP) (nam2020relative) considers bi-polar importance score in relevance by separating the relevant and irrelevant attributions. Relative Sectional Propagation (RSP) (nam2021interpreting) follows the propagation of the RAP rule while not suppressing the nature of activated neurons to enhance the explanation. Attribution Guided Factorization (AGF) (gur2021visualization) combines the accumulated gradients and relevance scores across the layers to correct the saliency bias.
@@
Despite the superior performance of the previous approach, the generated explanation maps represent insufficient saliency and diminished class discrimination. To increase the fidelity of the attribution maps, we consider different schemes of spatial scrutinization by providing multiple partial information from a single input image. Spatial scrutinization guides the attribution method to discover the obscured regions while maintaining class-discriminative saliencies.
@@
Attribution–based approaches provide a saliency map with the same size as the input image by propagating the relevance score with respect to the network decision. However, it is difficult to clarify the correlations of neurons among the whole layers, resulting in inevitably assigning the positive attributions in unrelated pixels. In other side, CAM–based approach provides the explanation based on the activation map of the deepest feature extraction layer and gradients, the computed saliency map represents a coarse explanation and lacks to details due to the upsampling. To provide more detail saliency maps, upscaling techniques of the input image is introduced in (jalwana2021cameras) to increasing the resolution of the class activation map.
(a) Deletion score
(b) Insertion score
Figure 2: The motivations of our work. The scores of deletion and insertion games among various resolutions with GradCAM on the ImageNet dataset. We scaled the size of image from $(224,224)$ to $(924,924)$ at intervals of $100$ pixels. Although two scores are in the trade-off relationships, the results tend to decrease in both cases.
@@
However, as shown in our empirical experiments of Figure 2, employing the upsampled images both decreases the deletion and insertion scores proportional to the resolution, which are originally in the trade-off relationships. Deletion and insertion scores are widely used metrics in the field of explanability (wang2020score; zhang2021novel; petsiuk2018rise; jung2021towards), to evaluate whether the explanation map successfully cover the more fine-grained representation and localizes the actual object better. Decreasing insertion scores according to the upscaling states that the visualized explanation loses saliency areas due to the distortion of neuron activation and corresponding gradients.
@@
Motivated from this issue, we address the method to improve the quality of saliency maps by investigating the correlations among the input unfolding and confidence scores, leading to maintaining the salient regions while increasing the granularity of the explanation map.
@@
Figure 3: An overview of UCAG. The entire course is divided in to three streams: i) unfoldment, ii) partial explanation, and iii) conquer with aggregation. Through the unfolding process, the unfolded patches are individually delivered to the gradient/attribution-based visualization methods. Conquering is a procedure for integrating the segmented explanations into a single explanation map with the corresponding confidence scores.
Unfold and Conquer Attribution Guidence
@@
To enhance the granularity of the explanation without losing saliencies, our work starts from the assumption that if the highlighted regions obviously contribute to the decision, their saliency should remain regardless of the geometrical translation. To fulfill this, we extend the aspect of convolution process generally used in the DNNs framework. Hence, we unfold the image into a series of segments and let the visualization method infers each fragment independently to yield partial explanation maps. By folding partial explanations weighted with corresponding confidence scores, the unified explanation contains the saliencies referring to the various perspectives of spatial views. Figure 3 illustrates the overview of our method: UCAG. Detailed description of each stage is described in each section.
@@
In this section, we demonstrate a strategy for providing the unfolded segments along the spatial axis to construct the set of input materials for the explanation method. To accomplish this, we first declare the hyper-parameter $n$: the number of segmented patches. Then, we construct a unfolding function: $\mathcal{U}$, which striding the input image $I\in\mathbb{R}^{C\times H\times W}$ into the series of $n^{2}$ patches $I_{p}$ by walking along the spatial (weight and height) axis. The set of patches includes the partial segment $I_{p}^{l}\in\mathbb{R}^{C\times H^{\prime}\times W^{\prime}}$, indexed by $l\in\{0,1,\dots,n^{2}-1\}$. The size of patch $(H^{\prime},W^{\prime})$ is computed with the specific resizing factor $\rho\in[0,1]$, described as $(\rho H,\rho W)$.
However, from the perspective of the network, utilizing the downscaled input leads to representing a coarse output features after passing through the network feature extraction. To relieve this, we enhance the resolution of each patch with the scaling function $\Psi(I_{p},\alpha)$, which scales the set of patches into the size of $(\alpha H^{\prime},\alpha W^{\prime})$. In sumamry, the unfoldment procedure provides the partial explanations as follows:
$I_{s}=\Psi(\mathcal{U}(I,\rho,n),\alpha).$ (1)
Hence, the input image $I$ is unfolded into the set of patches $I_{s}$ with size of $(n^{2},C,\alpha{H^{\prime}},\alpha{W^{\prime}})$ and each patch shares the overlapped areas with the striding mechanisms following the nature of unfoldment. From the spatial perspective, a set of patches holds different subregions from the single image and enables a network to review from the various views.
@@
Gradient–based saliency
To generate the saliency map revealing network decision, gradient–based methods utilize the activation map in the defined layers and their corresponding neuron importance, derived from the gradients. Generally, the activation map is obtained from the last layer of DNN before the linear or global average pooling layer, to investigate the compressed feature importance. Several CAM approaches introduce their intrinsic way of computing the class-specific coefficient by utilizing the gradient of each pixel in the activations. Generic gradient–based saliency map (selvaraju2017grad) is computed as follows:
@@
Here, $c$, $f$ and $k$ denote the index of the channel axis, the kind of DNNs, and index of the target class, respectively. $w_{c}^{k}$ is the neuron importance of each feature map derived from the gradients with respect to target class. According to the provided assumption of existing approaches, e.g. (selvaraju2017grad; fu2020axiom; chattopadhay2018grad; zhang2021novel), there is a slight difference of deriving the $w_{c}^{k}$, but mostly dependent on the chain rule ${\partial f(I)^{k}}/{\partial A}$. Detailed expansion of each method is described in supplementary materials. In our post-hoc manners: using the unfolded segments $I_{s}$ as input, partial explanations are generated independently with the activations computed from each patch as $\mathbb{E}_{grad}(I_{s},f,k)$.
Attribution–based saliency
Attribution–based approaches are capable of providing the input-level explanation map, allowing to represent pixel-level attribution. They compute the relevance score by layerwisly propagating the output logits to input in a backward manner. During the propagation, the whole relevance sum $R$ is maintained with the conservation rule(bach2015pixel), acting as the evidence of decision. The generic propagation rule (montavon2017explaining) is defined as follows:
@@
Here, $l$ indexes the middle layer of DNNs and $A$ represents the activation of corresponding layers. $i$ and $j$ denote the neuron index in the $l$ and $l-1$ layers, respectively. During the propagation, the whole relevance sum is maintained as $\sum_{j}R_{j}^{(l)}=\sum_{i}R_{j}^{(l-1)}$, starting from the output one-hot vector: $f(I)^{k}$ to the input layer.
Several kinds of attribution–based methods (nam2020relative; nam2021interpreting; bach2015pixel; gur2021visualization; gu2018understanding; iwana2019explaining; zhang2018top) derived from Eq.(3), but have the own techniques, e.g. uniform shifting and Factorization, to clearly visualize the final input attributions. Our method is independent of each of these unique techniques as $\mathbb{E}_{att}(I_{s},f,k)$ and could be utilized for the intuitive and clear visualizations.
Figure 4: Qualitative results of UCAG. UCAG improves the quality of the saliency map regardless of the type of model and method. Saliency maps marked with red boxes were originally generated from the attribution method, and those with blue boxes indicate that the maps were generated with UCAG.
@@
$\begin{split}Y_{s}^{k}&=\xi(\mathcal{L}_{s},k)\\ &=exp(softmax(\mathcal{L}_{s}^{k}).\end{split}$ (4)
$\mathcal{L}_{s}$ is the output logit of each segment: $f(I_{s})$ and $\xi(\cdot)$ returns the refined score with respect to the class $k$ by taking the exponential function to the softmax probability of $\mathcal{L}_{s}$, leading to expanding the gap of the partial confidence scores. Final explanation is aggregated by applying the element-wise product among the each partial explanation and the confidence scores. The weighted partial explanations is scaled to the original patch size: $(n^{2},1,H^{\prime},W^{\prime})$ by the scaling factor $\beta$ to return to the original areas of input image. In summary, the partial explanation before the folding function is computed as follows:
@@
Then, the scaling function $\Psi(\cdot)$ is applied to interpolate the series of explanations with the size of $(H^{\prime},W^{\prime})$. Note that the map $\mathbb{E}_{s}^{k}$ is not normalized or cliped (eliminating the negative scores) to gather more abundant range of explanations. Furthermore, since the partial explanations share the overlapped portions from the single image, the frequency of the pixels in the saliency map is not accumulated evenly, destabilizing the resulting saliency map. For this reason, we devise the duplication matrix $\Gamma$, which counts the duplication of each pixel. By utilizing this counts, we average each pixel in the saliency map according to their frequency of occurrence.
@@
The final saliency map has the same size of input image, representing the relative relevant scores in each pixel with respect to the output predictions of DNNs. The overall streams of our method is described in Alg.1.
@@
9:else if $\mathbb{E}$ is Attribution-based then
10:     $\mathbb{E}_{s}\leftarrow\mathbb{E}_{att}(I_{s},f,k)$triangleright Utilize attributions
@@
13:$Y_{k}\leftarrow\xi({L}_{s},k)$triangleright Compute confidence scores
14:$\mathbb{M}_{s}^{k}\leftarrow\Psi(\mathbb{E}_{s}^{k}\odot Y_{s}^{k},\beta)$triangleright Aggregation and Scaling
15:$\mathbb{M}^{k}\leftarrow\mathbb{V}(\mathbb{M}_{s}^{k},\Gamma_{s})$triangleright Average into final explanation
@@
We utilize the public accessible datasets: ImageNet-1k(deng2009imagenet), Pascal VOC 2002 (everingham2010pascal), and MS COCO (lin2014microsoft). For the localization and segmentation assessments, we additionally utilize the ImageNet segmentation dataset (guillaumin2014imagenet), which contains 4276 samples of images For the fair comparisons in evaluations, we employ the online available models: VGG19, DenseNet121, ResNet50, and InceptionV3.
We utilized several metrics: i) insertion and deletion(petsiuk2018rise), ii) (energy-based) pointing game, iii) mean Average Precison (mAP), iv) positive and negative density(jalwana2021cameras), to measure the degree of localization, causality, and density. Detailed comparison is described in each section. We set the hyperpameter for the number of segments and rescaling factor as $(\alpha=2.6,n=6,\rho=0.555)$. Comparisons according to the hyper-parameter are provided in the supplementary materials.
Figure 5: Qualitative results between multiple networks including ResNet50, DenseNet121, and InceptionV3. The top and bottom rows represent visualized heatmaps generated by GradCAM with and without our approach, respectively. More results are contained in the supplmentary material.

### v2

@@ Abstract
Revealing the transparency of Deep Neural Networks (DNNs) has been widely studied to describe the decision mechanisms of network inner structures. In this paper, we propose a novel post-hoc framework, Unfold and Conquer Attribution Guidance (UCAG), which enhances the explainability of the network decision by spatially scrutinizing the input features with respect to the model confidence. Addressing the phenomenon of missing detailed descriptions, UCAG sequentially complies with the confidence of slices of the image, leading to providing an abundant and clear interpretation. Therefore, it is possible to enhance the representation ability of explanation by preserving the detailed descriptions of assistant input features, which are commonly overwhelmed by the main meaningful regions. We conduct numerous evaluations to validate the performance in several metrics: i) deletion and insertion, ii) (energy-based) pointing games, and iii) positive and negative density maps. Experimental results, including qualitative comparisons, demonstrate that our method outperforms the existing methods with the nature of clear and detailed explanations and applicability.
@@ Introduction
Although Deep Neural Networks (DNNs) show remarkable performance in various computer science tasks due to their applicability, the issues of complexity and transparency derived from the non-linearity of the network hinder the clear interpretation of the decision basis. Visual explanations of network decisions alleviate these problems and provide analytical guidance for model development with intuitive explanations. Inspired by this, several approaches have been introduced, such as relevance or gradient-based generation, to describe the image regions that the network currently references to make decisions.
@@
Figure 1: The qualitative results of our approach. The first and second rows represent the existing visualization methods and ours, respectively. Our approach enhances the representation ability of explanation maps, which is applicable to any explanation technique.
@@
The approaches behind the relevance score generate the explanation based on the decomposition methods, Layer-wise Relevance Propagation (LRP) (Bach et al. 2015). LRP constructs the propagation rule according to the predicted scores of the specific class. Despite the nature of employing relevance score is robust to the shattered gradient problem and success in producing the input-level explanation map, the class-discriminativeness within their representation is still limited. Recently, although Attribution Guided Factorization (AGF) enhances the class discriminability of the explanation by integrating both gradient/relevance-based methodologies, it still provides biased attribution toward the class.
@@
GradCAM (Selvaraju et al. 2017) computes the coefficients for each map which are averaged over the gradients of the activation maps, and generates a class activation map by linearly combining the maps with the corresponding coefficients to provide a visual explanation. Although the explanation map generated from GradCAM successfully provides the class-discriminative explanation, the provided explanations are still coarse and contain numerous non-saliency regions due to the mechanism of utilizing the source map from the deepest layer, which keeps the lowest resolution activation map.
@@
The compressive nature of DNNs limits the representation ability of relevance/gradient-based approaches. As the features in the deep layer of the network hold the abbreviated information, it restricts the propagated regions to represent abundant saliencies. Thus, the resulting saliency maps produced from the gradient or relevance approach are represented with limited quality, with obscured class-discriminative regions. In this paper, we propose Unfold-and-Conquer Attribution Guidance (UCAG) framework to guide the explaining methods to generate better saliencies. It improves interpretability by investigating the detailed spatial regions with a set of subregions. Figure 1 illustrates the intuitive examples of our method, representing the strong advantages of detailed and abundant descriptions agnostic to models and methods. To fulfill this, UCAG is divided into three sections: spatial unfoldment, partial saliency generation, and conquer with geometrical aggregation. First, the unfoldment procedure provides a series of patches delivered by deploying the original data from a spatial perspective. The attribute method independently computes the partial explanation map. During this procedure, UCAG accepts any specific type of methodology as an explanation module, including explaining mechanisms and model types. Finally, the conquering procedure aggregates the partial saliency maps to produce the final explanatory map. Therefore, UCAG encourages an explanation method that scrutinizes an image through the spatial axis.
The key contributions of this paper are summarized as:
@@
  • We propose a new post-hoc framework: Unfold and Conquer Attribution Guidance (UCAG), to improve the interpretability of explanation methods by spatially scrutinizing the input data. Spatial scrutinization and integration provide multiple partial semantic views, leading to exploring the obscured saliencies due to the compressing nature of DNNs.
@@
  • Post-hoc manner of UCAG has the advantage of being easily applicable to various relevance/gradient-based approaches and DNN models. We present the applicability of UCAG that successfully provides a function of post-hoc refinement with intuitive and detailed descriptions of explanation.
@@
  • We validate the quality of the saliency map in density, localization, and causality perspectives. We conduct the verified experiments: i) positive and negative map density (Jalwana et al. 2021), ii) energy-based and original pointing game (Wang et al. 2020), and iii) insertion and deletion (Petsiuk, Das, and Saenko 2018) to confirm the perspectives, respectively. The results demonstrate that our method outperforms the existing explaining methods with a sizable gap in terms of indicating the better points in the input correlated with the output decision.
@@
Gradient-based approaches (Simonyan, Vedaldi, and Zisserman 2013; Zeiler and Fergus 2014; Springenberg et al. 2014; Selvaraju et al. 2017; Chattopadhay et al. 2018; Srinivas and Fleuret 2019; Rebuffi et al. 2019; Fu et al. 2020; Rebuffi et al. 2020; Zhang, Rao, and Yang 2021) construct the saliency map derived from the gradients. In the manner of employing class activation maps, which employs the activation map to generate the saliency map, GradCAM (Selvaraju et al. 2017) alleviates the limitations of CAM that require a global average pooling and re-training process. They construct an explanatory map by linearly combining the activation map and the coefficients with the coefficients computed by taking the average of the gradients corresponding to the activation map. XGradCAM (Fu et al. 2020) introduces the way to generate the class activation map while obeying two theoretical axioms: sensitivity and conservation. WGradCAM (Zhang, Rao, and Yang 2021) enhances the representation ability of the explanation map in terms of localization.
@@
Several approaches (Fong, Patrick, and Vedaldi 2019; Jiang et al. 2021; Jalwana et al. 2021; Zhang, Rao, and Yang 2021) turned the literature to amplifying the quality of saliency maps to produce a more precise saliency map. Extremal perturbation (Fong, Patrick, and Vedaldi 2019) generates the saliency map following the searching nature of perturbation-based approaches. LayerCAM (Jiang et al. 2021) aggregates the information from earlier layers and introduces the way of deciding the coefficients of each source. CAMERAS (Jalwana et al. 2021) fuses the multi-resolution images to generate a high-resolution explanation while obeying the model fidelity. SigCAM (Zhang, Rao, and Yang 2021) combines perturbation/gradient-based approaches to produce more sparse and precise explanations using learnable external parameters.
Relevance-based approaches (Bach et al. 2015; Zhang et al. 2018; Gu, Yang, and Tresp 2018; Iwana, Kuroki, and Uchida 2019; Nam et al. 2020; Nam, Choi, and Lee 2021; Gur, Ali, and Wolf 2021) produce the granular explanation by decomposing the output predictions in a backward manner. LRP (Bach et al. 2015) introduces the propagation rules that decompose the output logits into the relevance scores while maintaining the conservation rule. As the recent research, RAP (Nam et al. 2020) aims to separate the relevant and irrelevant parts of input by allocating the bi-polar relevance scores during propagation. RSP (Nam, Choi, and Lee 2021) extends the rule of RAP to overcome the limitations of suppressing nature, resulting in solving the class-specific issues. AGF (Gur, Ali, and Wolf 2021) combines the accumulated gradients and relevance scores across the layers to correct the saliency bias.
@@
Despite the superior performance of the previous approach, the generated explanation maps represent insufficient saliency and diminished class discrimination. To increase the fidelity of the explanation maps, we consider the schemes of spatial scrutinization and integration by providing multiple partial information from a single input image, leading to discovering the obscured regions while maintaining class-discriminative saliencies.
@@
Propagation-based approaches provide a saliency map with the same size as the input image by propagating the relevance score with respect to the network decision. However, it is difficult to clarify the correlations of neurons among the whole layers, resulting in inevitably assigning positive attributions to unrelated pixels. On another side, since the CAM-based approach provides the explanation based on the activation map of the deepest feature extraction layer and gradients, the computed saliency map represents a coarse explanation and lacks details due to the upsampling. To provide more detailed saliency maps, upscaling techniques of the input image are introduced in (Jalwana et al. 2021) to increase the resolution of the class activation map.
Figure 2: The motivations of our work. The scores of deletion and insertion games among various resolutions with GradCAM on the ImageNet dataset. We scaled the size of image from $(224,224)$ to $(1120,1120)$ utilizing the scale factor $\alpha$. The performance of the original and our methods are represented as dots and solid lines, respectively. We aim to present a resolution-agnostic method with the enhancement of interpretability.
@@
However, as shown in our empirical experiments of Figure 2, employing the upsampled images both decreases the deletion and insertion scores proportional to the resolution, which are originally in the trade-off relationships. Deletion and insertion scores are widely used metrics in the field of explainability  (Petsiuk, Das, and Saenko 2018; Wang et al. 2020; Zhang, Rao, and Yang 2021; Jung and Oh 2021), to evaluate whether the explanation map successfully covers the more fine-grained representation and localizes the actual object better. Decreasing insertion scores according to the upscaling states that the visualized explanation loses saliency areas due to the distortion of neuron activation and corresponding gradients.
@@
Motivated by this issue, we address the method to improve the quality of saliency maps by investigating the correlations among the input unfolding and confidence scores, leading to maintaining the salient regions while increasing the granularity of the explanation map.
@@
Figure 3: An overview of UCAG. The entire course is divided into three streams: i) unfoldment, ii) patch-wise saliency generation, and iii) conquer with aggregation. Through the unfolding process, the unfolded patches are individually delivered to the gradient/propagation-based visualization methods. Conquering is a procedure for integrating segmented explanations into a single explanation map with the corresponding confidence scores.
Unfold and Conquer Attribution Guidance
@@
To enhance the granularity of the explanation without losing saliencies, our work starts from the assumption that if the highlighted regions obviously contribute to the decision, their saliency should remain regardless of the geometrical translation. To fulfill this, we extend the aspect of the convolution process generally used in the DNNs framework. Hence, we unfold the image into a series of segments and let the visualization method infers each fragment independently to yield partial explanation maps. By folding partial explanations weighted with corresponding confidence scores, the unified explanation contains the saliencies referring to the various perspectives of spatial views. Figure 3 illustrates the overview of our method: UCAG. A detailed description of each stage is described in each section.
@@
In this section, we demonstrate a strategy for providing the unfolded segments along the spatial axis to construct the set of input materials for the explanation method. To accomplish this, we first declare the hyper-parameter $n$: the number of segmented patches. Then, we construct an unfolding function: $\mathbb{U}$, which striding the input image $I\in\mathbb{R}^{C\times H\times W}$ into the series of $n^{2}$ patches $I_{p}$ by walking along the spatial (weight and height) axis. The set of patches includes the partial image $I_{p}^{j}\in\mathbb{R}^{C\times H^{\prime}\times W^{\prime}}$, indexed by $j\in\{0,1,\dots,n^{2}-1\}$. The size of patch $(H^{\prime},W^{\prime})$ is computed with the specific resizing factor $\rho\in[0,1]$, described as $(\rho H,\rho W)$.
However, from the perspective of the network, utilizing the downscaled input leads to representing coarse output features after passing through the network feature extraction. To relieve this, we enhance the resolution of each patch with the scaling function $\Psi(I_{p},\alpha)$, which scales the set of patches into the size of $(\alpha H^{\prime},\alpha W^{\prime})$. In summary, the unfoldment procedure provides partial explanations as follows:
$I_{s}=\Psi(\mathbb{U}(I,\rho,n),\alpha).$ (1)
Hence, the input image $I$ is unfolded into the set of patches $I_{s}$ with a size of $(n^{2},C,\alpha{H^{\prime}},\alpha{W^{\prime}})$ and each patch shares the overlapped areas with the striding mechanisms following the nature of unfoldment. From the spatial perspective, a set of patches holds different subregions from the single image and enables a network to review the various views.
@@
Gradient-based saliency
To generate the saliency map revealing network decisions, gradient-based methods utilize the activation map in the defined layers and their corresponding neuron importance derived from the gradients. Generally, the activation map is obtained from the last layer of DNN before the linear or global average pooling layer to investigate the compressed feature importance. Several CAM approaches introduce their intrinsic way of computing the class-specific coefficient by utilizing the gradient of each pixel in the activations. Generic gradient-based saliency map (Selvaraju et al. 2017) is computed as follows:
@@
Here, $c$, $f$, and $k$ denote the index of the channel axis, the kind of DNNs, and the index of the target class, respectively. $w_{c}^{k}$ is the neuron importance of each feature map derived from the gradients with respect to the target class. According to the provided assumption of existing approaches, e.g. (Selvaraju et al. 2017; Chattopadhay et al. 2018; Fu et al. 2020; Zhang, Rao, and Yang 2021), there is a slight difference in deriving the $w_{c}^{k}$, but mostly dependent on the chain rule ${\partial f(I)^{k}}/{\partial A}$. Detailed expansion of each method is described in supplementary materials. In our post-hoc manners: using the unfolded segments $I_{s}$ as input, partial explanations are generated independently with the activations computed from each patch as $\mathbb{E}_{grad}(I_{s},f,k)$.
Propagation-based saliency
Propagation-based methods are capable of providing the input-level explanation map, allowing for the representation of pixel-level attribution. They compute the relevance score by layer-wisely propagating the output logits to input in a backward manner. During the propagation, the whole relevance sum $R$ is maintained with the conservation rule (Bach et al. 2015), acting as the evidence of the decision. The generic propagation rule (Montavon et al. 2017) is defined as follows:
@@
Here, $l$ indexes the middle layer of DNNs, and $A$ represents the activation of corresponding layers. $i$ and $j$ denote the neuron index in the $l$ and $l-1$ layers, respectively. During the propagation, the whole relevance sum is maintained as $\sum_{j}R_{j}^{(l)}=\sum_{i}R_{j}^{(l-1)}$, starting from the output one-hot vector: $f(I)^{k}$ to the input layer.
Several kinds of propagation-based methods, e.g., SGLRP, c*LRP, RSP, and AGF are mainly based on Eq. 3, but have some techniques, such as contrastive perspective, uniform shifting, and factorization, to enhance the quality of attribution maps. Our method is independent of each of these unique techniques as $\mathbb{E}_{att}(I_{s},f,k)$ and could be utilized for intuitive and clear visualizations.
Figure 4: Qualitative comparisons in VGG-19 and ResNet-50. UCAG improves the quality of the saliency map and has an advantage of agnostic to model and method. First and second rows in each group represent the original and our (marked as red) results, respectively. Additional results are illustrated in supplementary materials.
@@
$\begin{split}Y^{k}_{s}&=\xi(\mathcal{L}_{s},k)\\ &=\textrm{exp}\left(\frac{\textrm{exp}(\mathcal{L}_{s}^{k,j})}{\sum_{i=1}^{N}\textrm{exp}(\mathcal{L}_{s}^{i,j})}\right),\forall j.\end{split}$ (4)
Here, $j\in[0,n^{2}-1]$ indexes the logit, $\mathcal{L}_{s}$ is the output logit: $f(I_{s})$ and $\xi(\cdot)$ returns the refined score with respect to the class $k$ by individually taking the exponential function to the softmax probability of $\mathcal{L}^{j}_{s}$, leading to expanding the gap of the partial confidence scores. $N$ represents the number of classes. The final explanation is aggregated by applying the element-wise product among each partial explanation and the confidence scores. The weighted partial explanations are scaled to the original patch size: $(n^{2},1,H^{\prime},W^{\prime})$ by the scaling factor $\beta$ to return to the original areas of the input image. In summary, the partial explanation before the folding function is computed as follows:
@@
Then, the scaling function $\Psi(\cdot)$ is applied to interpolate the series of explanations with the size of $(H^{\prime},W^{\prime})$. Note that the map $\mathbb{E}_{s}^{k}$ is not normalized or clipped (eliminating the negative scores) to gather a more abundant range of explanations. Furthermore, since the partial explanations share the overlapped portions from the single image, the frequency of the pixels in the saliency map is not accumulated evenly, destabilizing the resulting saliency map. For this reason, we devise the duplication matrix $\Gamma$, which counts the duplication of each pixel. By utilizing this count, we average each pixel in the saliency map according to its frequency of occurrence.
@@
The final saliency map has the same size as the input image, representing the relatively relevant scores in each pixel with respect to the output predictions of DNNs. The overall streams of our method are described in Algorithm.1.
@@
9:else if $\mathbb{E}$ is Propagation-based then
10:     $\mathbb{E}_{s}\leftarrow\mathbb{E}_{att}(I_{s},f,k)$triangleright Utilize propagations
@@
13:$Y^{k}_{s}\leftarrow\xi({L}_{s},k)$triangleright Compute confidence scores
14:$\mathbb{M}_{s}^{k}\leftarrow\Psi(\mathbb{E}_{s}^{k}\odot Y^{k}_{s},\beta)$triangleright Aggregation and Scaling
15:$\mathbb{M}^{k}\leftarrow\mathbb{V}(\mathbb{M}_{s}^{k},\Gamma_{s})$triangleright Generating final explanation
@@
We utilize the public accessible datasets: ImageNet-1k (Deng et al. 2009), Pascal VOC 2002 (Everingham et al. 2010), and MS COCO (Lin et al. 2014). For the localization and segmentation assessments, we utilize the ImageNet segmentation dataset (Guillaumin, Küttel, and Ferrari 2014). For fair comparisons in evaluations, we employ the online available models: VGG19 (Simonyan and Zisserman 2014), DenseNet121 (Huang et al. 2017), ResNet50 (He et al. 2016), and InceptionV3 (Szegedy et al. 2016).
We utilized several metrics: i) insertion and deletion games, ii) (energy-based) pointing game, iii) mean Average Precision (mAP), and iv) positive and negative map density to measure the degree of localization, causality, and density. A detailed comparison is described in each section. We set the hyperparameter for the number of segments and rescaling factor as $(\alpha=2.6,n=6,\rho=0.555)$. Comparisons according to the hyper-parameter are provided in the supplementary materials.
Figure 5: Qualitative results between multiple networks including ResNet50, DenseNet121, and InceptionV3. The top and bottom rows represent visualized heatmaps generated by GradCAM with and without our approach, respectively. More results are contained in the supplementary material.